SPECTROSCOPIC AND KINETIC MEASUREMENTS OF ORGANIC PEROXY RADICALS BY DUAL-WAEVELENGTH CAVITY RING DOWN SPECTROSCOPY

\underline{DMITRY G. MELNIK}, and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

The weakly-absorbing Ã ← X̃ transition of organic peroxy radicals in the NIR provides an excellent opportunity for quantitative species-specific measurements. Previously^{a} we used the dual-wavelength cavity ring-down spectroscopy (CRDS) technique to accurately measure the absolute peak absorption cross-section of ethyl peroxy radical, C_{2}H_{5}O_{2}. In the present work, we extend the capabilities of the dual-wavelength technique to studying kinetics of the self-reaction of ethyl peroxy by utilizing a high duty-factor, time-resolved CW-CRDS. An extension of this technique to study cross-reactions between different species will be discussed.

aD. Melnik et al., J. Phys. Chem. A, 114, 11583 (2010)